### v1

# The dual complex of a $G$-variety

Abstract.
We introduce a new invariant of $G$-varieties, the dual complex, which roughly measures how divisors in the complement of the free locus intersect. We show that the top homology group of this complex is an equivariant birational invariant of $G$-varieties. As an application, we demonstrate the non-linearizability of certain large abelian group actions on smooth hypersurfaces in projective space of any dimension and degree at least $3$.
Key words and phrases:
equivariant birational geometry, hypersurfaces
2020 Mathematics Subject Classification:
14L30, 14E08 (primary); 14J70 (secondary)
1. Introduction
Let $G$ be a finite group which acts on a smooth projective variety $X$. A major problem in algebraic geometry is the classification of such $G$-varieties up to $G$-birational equivalence, where $X$ and $Y$ are $G$-birationally equivalent if there are invariant open sets $U\subset X$, $V\subset Y$ and an equivariant isomorphism $\varphi:U\xrightarrow{\cong}V$.
The geometry of $G$-varieties may be packaged into invariants that distinguish $G$-birational equivalence classes. This circle of ideas begins with the observation that the presence of an $H$-fixed point is a $G$-birational invariant of $X$ for $H\subset G$ abelian, over an algebraically closed field of characteristic zero [15, App. A]. The recent introduction of equivariant Burnside groups [11] has led to many new advances in this area. The $G$-birational invariants which take values in these groups incorporate the birational types of fixed loci of all abelian subgroups $H\subset G$, actions of $H$ on the normal spaces to these fixed loci, and residual actions on their orbits. They have found numerous applications in classifying $G$-varieties of low dimension, such as projective representations of finite groups [12, 18], algebraic tori [13], and actions on other rationally connected threefolds [5].
Another tool with significant applications in birational geometry is the dual complex of a simple normal crossings divisor. This is a simplicial (quasi-)complex which describes how the irreducible components of the divisor intersect. Such objects may be associated to (log resolutions of) singularities [6] and Calabi-Yau pairs [10] in such a way that the homotopy type of the complex is independent of the resolution chosen.
Existing $G$-birational invariants do not directly measure how components of the non-free locus of a $G$-variety intersect. Therefore, in this paper, we introduce a notion of dual complexes for $G$-varieties, which record intersection data for the part of the non-free locus that “looks most like the boundary of a toric variety”. We show that the top homology group of this complex is a $G$-birational invariant (Theorem 2.4) and use this result to find new examples of non-birational actions. In particular, we prove that actions of rank $n$ abelian subgroups $G\subset\mathrm{Aut}(\mathbb{P}^{n+1})$ on invariant smooth hypersurfaces $X\subset\mathbb{P}^{n+1}$ of degree $d\geq 3$ are never $G$-birational to a toric action on a smooth toric variety, except possibly in the case of cubic surfaces (Theorem 3.2).
1.1. Notation
$G$ will always denote a finite group. We will only consider smooth and projective $G$-varieties $X$ with faithful action. All varieties are over an algebraically closed field $k$ of characteristic zero. For a subgroup $H\subset G$, $X^{H}$ is the locus in $X$ on which $H$ acts trivially. We say that a $G$-variety has abelian stabilizers if for every closed point $x\in X$, $\mathrm{stab}_{G}(x)\coloneqq\{g\in G:g\cdot x=x\}$ is abelian. The rank of a finite abelian group $H$ is the minimal number of generators of $H$.
Acknowledgements. Thank you to Nathan Chen, János Kollár, Burt Totaro, Ravi Vakil, and Ziquan Zhuang for useful conversations and comments. Thank you to Yuri Tschinkel for suggesting an alternative approach to Theorem 3.2.
2. Dual Complexes
Suppose $X$ is a smooth projective $G$-variety with abelian stabilizers. By Luna’s slice theorem [14, Lemme III.1], for any closed point $p\in X$, there is an invariant affine open neighborhood $U$ of $p$ and a $\mathrm{stab}_{G}(p)$-equivariant morphism $\varphi:U\rightarrow T_{p}X$ such that $\varphi$ is étale at $p$ and $\varphi(p)=0$. In other words, the action is “étale-locally linear”. Therefore, for any subgroup $H\subset\mathrm{stab}_{G}(p)$, $T_{p}X^{H}\cong(T_{p}X)^{H}$ and $X^{H}$ is smooth at $p$ (though globally $X^{H}$ may be a union of connected components of different dimensions). It follows that the representation of $\mathrm{stab}_{G}(p)$ on $T_{p}X$ is faithful. Thus, for any abelian group $H\subset G$, any connected component of $X^{H}$ has codimension at least $\mathrm{rank}(H)$ in $X$.
The connected components $Z$ of $X^{H}$ for which ${\operatorname{codim}}(Z)=\mathrm{rank}(H)$ are particularly well-behaved, by the following lemma.
Lemma 2.1.
Let $X$ be a $G$-variety with abelian stabilizers, $H\subset G$ an abelian subgroup, and $Z$ a connected component of $X^{H}$ such that ${\operatorname{codim}}(Z)=\mathrm{rank}(H)=k$. Then $Z$ is a component of an intersection $D_{1}\cap\cdots\cap D_{k}$ of divisors, where the general point $x\in D_{i}$ has nontrivial cyclic stabilizer $H_{i}$, and $H=\bigoplus_{i=1}^{k}H_{i}$. Moreover, these divisors have simple normal crossings in a neighborhood of $Z$.
Proof.
For any point $p\in Z$, we may diagonalize the action of $H$ on $T_{p}X$ to $T_{p}Z\oplus\chi_{1}\oplus\cdots\oplus\chi_{k}$, where the $\chi_{i}$ are characters of $H$ that do not depend on the choice of point $p\in Z$. The representation $\chi_{1}\oplus\cdots\oplus\chi_{k}$ of $H$ must be faithful. Since $H$ has rank $k$, it follows that $H=\bigoplus_{i=1}^{k}H_{i}$, where
$H_{i}\coloneqq\ker(\chi_{1})\cap\cdots\cap\widehat{\ker(\chi_{i})}\cap\cdots\cap\ker(\chi_{k})$
is a nontrivial cyclic group for each $i$. The subgroup $H_{i}$ is the stabilizer of a hyperplane in $T_{p}X$, so this hyperplane is the tangent space to a divisor $D_{i}$ in $X^{H_{i}}$ passing through $p$. The divisors $D_{1},\ldots,D_{k}$ have simple normal crossings at $p$ since all of their intersections are smooth of the expected dimension near $p$ [17, Lemma 0BIA]. ∎
We refer to the components $Z$ of $X^{H}$ with ${\operatorname{codim}}(Z)=\mathrm{rank}(H)$ as the maximal rank strata of the $G$-variety $X$. Next, we construct a complex measuring the intersections of these special strata.
Definition 2.2.
The dual complex $\mathcal{D}_{G}(X)$ of a smooth projective $G$-variety $X$ of dimension $n$ is a CW complex constructed as follows. The 00-skeleton of $\mathcal{D}_{G}(X)$ is a union of points indexed by divisorial components of $X^{H}$, as $H$ ranges over all nontrivial cyclic abelian subgroups of $G$. At the $k$th step, attach a $(k-1)$-simplex for each connected component $Z$ of $X^{H}$ with ${\operatorname{codim}}(Z)=\mathrm{rank}(H)$, for $H\subset G$ abelian of rank $k$. By Lemma 2.1, $Z$ is a connected component of the intersection $D_{1}\cap\cdots\cap D_{k}$, where the divisors $D_{1},\ldots,D_{k}$ each have nontrivial stabilizer and simple normal crossings along $Z$. The new simplex is attached along the simplices corresponding to $D_{1},\ldots,D_{k}$ and the components of their intersections containing $Z$.
The result of this construction is a simplicial quasicomplex $\mathcal{D}_{G}(X)$ of dimension at most $n-1$. That is, it is a CW complex whose cells are simplices such that the collection of simplices is closed under the face relation and the intersection of any two is a union of some of their faces. We refer to [2, Section 7] for more about quasicomplexes and related terminology. {commentA} We next review a some notions for quasicomplexes which naturally arise when considering the dual complexes of blowups.
Definition 2.3.
Let $\Sigma$ be a simplicial quasicomplex. We use the following terminology:
  • The star $\mathrm{Star}(\Sigma,\Delta)$ for $\Delta\in\Sigma$ is the set of all faces of $\Sigma$ containing $\Delta$
  • The closure $\bar{S}\subset\Sigma$ of a subcollection $S\subset\Sigma$ consists of every simplex in $S$ and its faces
  • The link of $\Delta$ is $L(\Sigma,\Delta)\coloneqq\overline{\mathrm{Star}(\Sigma,\Delta)}\setminus\mathrm{Star}(\Sigma,\Delta)$
  • The stellar subdivision of a complex $\Sigma$ in a simplex $\Delta$ with barycenter $p$ is the simplicial subcomplex
    $\Sigma\setminus\mathrm{Star}(\Sigma,\Delta)\cup\{\mathrm{conv}(p,\Delta)|\Delta\in L(\Sigma,\Delta)\}$
Theorem 2.4.
The homology group $H_{n-1}(\mathcal{D}_{G}(X))$ is a $G$-birational invariant of smooth projective varieties of dimension $n$ with abelian stabilizers.
Proof.
By the $G$-equivariant weak factorization theorem [1], any birational map of smooth projective $G$-varieties may be decomposed into a sequence of blowups and blowdowns in smooth centers. If the varieties both have abelian stabilizers, we may guarantee that the intermediate varieties have abelian stabilizers as well (cf. [11, Proposition 3.6]). {commentA} Indeed, that proposition shows that any two varieties $X$ and $X^{\prime}$ satisfying Assumption 2 of [11] are connected by a series of blowups and blowdowns in smooth centers, where each intermediate variety satisfies Assumption 2 as well. One of the conditions of Assumption 2 is having abelian stabilizers. Though this isn’t sufficient to satisfy Assumption 2, there are equivariant blowups of $X$ and $X^{\prime}$ satisfying Assumption 2 [11, Proposition 3.3], and these can be connected by intermediate varieties satisfying the assumption. Since a blowup of a variety with abelian stabilizers has abelian stabilizers, this completes the argument. Therefore, it suffices to show that the homology group $H_{n-1}$ of the dual complex is unchanged by a single blowup $X^{\prime}\coloneqq\mathrm{Bl}_{W}(X)\rightarrow X$ of a variety $X$ of dimension $n$ with abelian stabilizers in a $G$-stable center $W$, which is a disjoint union of smooth irreducible varieties. We may assume that $G$ acts transitively on the connected components of $W$.
There are two cases to consider. First, suppose that the rank of the stabilizer of a general point in $W$ equals ${\operatorname{codim}}(W)$. Then each component of $W$ is a maximal rank stratum. We claim that in this case $\mathcal{D}_{G}(X^{\prime})$ is the stellar subdivision of the complex $\mathcal{D}_{G}(X)$ in the simplices corresponding to the components of $W$. Note that distinct components of $W$ are disjoint, so no simplex in $\mathcal{D}_{G}(X)$ contains multiple faces corresponding to components of $W$. It follows that we can perform all these stellar subdivisions simultaneously, and it’s enough to consider the neighborhood of a single component.
To see that the effect of the blowup is a stellar subdivision, let $Z$ be any connected component of $W$, and $D_{1},\ldots,D_{k}$ the divisors from Lemma 2.1 containing $Z$. Each $D_{i}$ is fixed by a nontrivial cyclic subgroup $H_{i}\subset G$ and the sum of these subgroups has rank $k$. Denote by $E_{Z}$ the exceptional divisor over $Z$. Any maximal rank stratum contained in $Z$ is a component of an intersection of the form $D_{1}\cap\cdots\cap D_{k}\cap D_{1}^{\prime}\cap\cdots\cap D_{\ell}^{\prime}$ for some additional divisors $D_{1}^{\prime},\ldots,D_{\ell}^{\prime}$, where the stabilizer of a general point on each is a nontrivial cyclic group. Furthermore, these divisors have simple normal crossings in a neighborhood of the intersection. It’s well known that the dual complex of the blow up of a stratum in an snc intersection is the stellar subdivision of the original complex (see, e.g., [2, Section 7]). It remains to see that all the strata of the blowup “count” for our notion of dual complex, namely that for each proper subset $I\subset\{1,\ldots,k\}$, $E_{Z}\cap D_{I}\cap D_{1}^{\prime}\cap\cdots\cap D_{\ell}^{\prime}$ is a maximal rank stratum in $X^{\prime}$, where $D_{I}=\bigcap_{i\in I}D_{i}$.
Indeed, the stabilizer of a general point $p$ in $E_{Z}$ is the intersection of $H\subset\mathrm{GL}((N_{Z/X})_{p})$ with the scalar transformations on this normal space, which must be a nontrivial cyclic group since $\mathrm{rank}(H)=\dim((N_{Z/X})_{p})$. This group and any $|I|<k$ summands $H_{i}$ of $H$ generate a rank $|I|+1$ abelian subgroup. The intersection $E_{Z}\cap D_{I}\cap D_{1}^{\prime}\cap\cdots\cap D_{\ell}^{\prime}$ is therefore fixed by a group of rank $|I|+1+\ell$, as desired. Since maximal rank strata are intersections of maximal rank divisors, no other new maximal rank strata are introduced by the blowup. Stellar subdivision induces a homeomorphism of complexes, so the homology of $\mathcal{D}_{G}(X^{\prime})$ is the same as that of $\mathcal{D}_{G}(X)$ in this case.
The second case is that the stabilizer of a general point in $W$ has rank smaller than ${\operatorname{codim}}(W)$. For this case, we require the following lemma:
Lemma 2.5.
Let $H={\mathbb{Z}}/d_{1}\oplus\cdots\oplus{\mathbb{Z}}/d_{n}$ be a rank $n$ abelian group acting diagonally on $\mathbb{A}^{n}$. Then any $H$-stable smooth subvariety $Y\subset\mathbb{A}^{n}$ through $0\in\mathbb{A}^{n}$ is a coordinate plane.
Proof.
Let $Y=V(I)$ for $I\subset k[x_{1},\ldots,x_{n}]$ a prime ideal. Since $I$ is $H$-invariant, we may choose a generating set $S\coloneqq\{f_{1},\ldots,f_{t}\}$ for $I$ such that $g\cdot f_{i}=\chi_{i}(g)f_{i}$ for some character $\chi_{i}$ of $H$.
{commentA}
To prove this, let $T\coloneqq\{h_{1},\ldots,h_{s}\}$ be any generating set. We can create a finite $H$-invariant generating set $S$ of $I$ by taking the $H$-orbits of all elements of $T$. Then $\mathrm{span}_{k}(S)$ is a finite-dimensional vector space with $H$-action. We can diagonalize the action, so that each basis element is preserved by $H$ up to scaling. This basis is also a generating set for $I$ with the required property.
Since $Y$ is smooth and passes through 00, its tangent cone is some linear subspace of $\mathbb{A}^{n}$. It is also $H$-stable, so we may assume the initial ideal of $I$ (which cuts out the tangent cone) is $\mathrm{in}(I)=(x_{1},\ldots,x_{r})$, where $\dim(Y)=n-r$. Among the linear forms in $\mathrm{in}(I)$, only $x_{1},\ldots,x_{r}$ are preserved up to scaling by $H$, so the generating set contains an element of the form $x_{1}+h_{1}$, where $h_{1}$ only has terms of degree at least $2$. Since $x_{1}$ and $h_{1}$ must have the same $G$-degree, it follows that $x_{1}$ divides $h_{1}$. Since $I$ is prime, $x_{1}\in I$. Similarly, $x_{2},\ldots,x_{r}\in I$. For dimension reasons, $I=(x_{1},\ldots,x_{r})$, so $Y$ is a coordinate plane. ∎
Returning to the proof of Theorem 2.4, we claim that the assumption on the stabilizers of points of $W$ means that $W$ does not contain a maximal rank stratum of dimension 00. Suppose by way of contradiction that $W$ does contain such a point $p$ with stabilizer $H$, with $\mathrm{rank}(H)=n$. There is a neighborhood $U$ of $p$ and an $H$-equivariant morphism $\varphi:U\rightarrow T_{p}X$ which is étale at $p$ and sends $p$ to 00, by Luna’s slice theorem. Restricting to a smaller open $V\subset U$ containing $p$ where $\varphi$ is étale, we have that $\phi(V\cap W)$ is an $H$-stable closed subvariety of an open neighborhood of 00 in $T_{p}X\cong\mathbb{A}^{n}$ which passes through 00 and is smooth there. The group $H$ is rank $n$ and acts linearly on $T_{p}X$. After diagonalizing this action, we may apply Lemma 2.5 to conclude that $\varphi(V\cap W)$ coincides with a coordinate plane near 00 in $T_{p}X$. Every coordinate plane is the image of a maximal rank stratum, so $W$ and some stratum $Z$ coincide étale-locally near $x$. Hence $W$ is a union of maximal rank strata, a contradiction.
Since $W$ contains no point with stabilizer of rank $n$, it also does not contain any stratum containing such a point. Therefore, the subcomplex of $\mathcal{D}_{G}(X)$ consisting of the closures of all $(n-1)$-simplices remains the same in $\mathcal{D}_{G}(X^{\prime})$, where each stratum is replaced by its strict transform, and a collection of strata intersects in $X$ if and only if it does in $X^{\prime}$. There could be new cells in $\mathcal{D}_{G}(X^{\prime})$ corresponding to exceptional divisors over $W$ and their intersections with existing strata. However, since every point of $W$ has stabilizer of rank less than $n$, all new $k$-cells satisfy $k<n-1$ and don’t appear in the boundary of an $(n-1)$-simplex. Therefore, the boundary map on cells of dimension $n-1$ remains unchanged, and the homology groups $H_{n-1}(\mathcal{D}_{G}(X))$ and $H_{n-1}(\mathcal{D}_{G}(X^{\prime}))$ are isomorphic. ∎
The other homology groups of the dual complex are not $G$-birational invariants. As a simple example, suppose the action of $G={\mathbb{Z}}/2$ on a smooth variety $X$ of dimension at least $2$ has an isolated fixed point $p$. Since the codimension of $p$ is greater than $1$, the point is not a maximal rank stratum and is not accounted for in the dual complex. However, blowing up $p$ gives a new fixed divisor, which will increase the number of connected components of the dual complex by one.
3. Applications
In this section, we consider applications of the invariant $H_{n-1}(\mathcal{D}_{G}(X))$ defined in Section 2 to the classification of $G$-actions up to birational equivalence. First, note that this group always vanishes unless the dual complex contains $(n-1)$-cells, which can only occur if $G$ contains an abelian group of rank at least $n=\dim(X)$. Toric varieties form one natural class with actions of this kind.
Example 3.1 (Smooth Toric Varieties).
Let $X$ be a smooth projective toric variety of dimension $n$ and $G\subset\mathbb{G}_{\operatorname{m}}^{n}$ a full rank finite subgroup of the torus. The non-free locus for the $G$-action is precisely the toric boundary, and each stratum of the boundary has maximal rank by the assumption that $G$ has rank $n$. {commentA} For instance, consider any toric divisor of $X$, corresponding to a ray with primitive point $r=(c_{1},\ldots,c_{n})$ in the lattice $N$ of one-parameter subgroups of $\mathbb{G}_{\operatorname{m}}^{n}$. Our claim amounts to the fact that any such one-parameter subgroup intersects $G$ nontrivially. The group fixing $D_{r}$ for $r$ as above is $\{(t^{c_{1}},\ldots,t^{c_{n}}):t\in{\mathbb{C}}^{*}\}$. Suppose $G$ has elementary abelian subgroup $({\mathbb{Z}}/p)^{\oplus n}$. Then plugging in $t=\zeta_{p}$, we find a nontrivial element $(\zeta_{p}^{c_{1}},\ldots,\zeta_{p}^{c_{n}})$ in the intersection of $G$ with the one-parameter subgroup (it is nontrivial because not all $c_{i}$ can be divisible by $p$). The rays spanning any cone in the fan of $X$ form part of a ${\mathbb{Z}}$-basis for $N$, so we can use a similar argument to find a subgroup of the appropriate rank fixing the corresponding troric stratum. Since $X$ is smooth and projective, its fan is the span of the faces of some simplicial polytope $P$ in the lattice $N$ of one-parameter subgroups [9, Theorem VII.3.11]. Therefore, the dual complex $\mathcal{D}_{G}(X)$ is simply $P$. In particular, $\mathcal{D}_{G}(X)$ is homeomorphic to the sphere $S^{n-1}$, so the top homology of $\mathcal{D}_{G}(X)$ is $H_{n-1}(S^{n-1})\cong{\mathbb{Z}}$.
A faithful $G$-action on a variety $X$ is linearizable if there is a $G$-birational map $X\dashrightarrow\mathbb{P}(V)$, where the action on $\mathbb{P}(V)$ is via a faithful linear representation $G\rightarrow\mathrm{GL}(V)$. When $G$ is abelian, we may diagonalize the $G$-action on $V$ to show that $G$ is a subgroup of the torus in $\mathbb{P}(V)$. If $G$ also has rank $n=\dim(X)$, then Theorem 2.4 and Example 3.1 show that $H_{n-1}(\mathcal{D}_{G}(X))\cong{\mathbb{Z}}$ whenever the action on $X$ is linearizable.
Rank $n$ abelian subgroups also act on many hypersurfaces in $\mathbb{P}^{n+1}$. We demonstrate that these actions are never linearizable when the degree of the hypersurface is at least $3$, except for the case of cubic surfaces.
Theorem 3.2.
Let $n$ and $d$ be positive integers such that $n\geq 1$, $d\geq 3$, and $(n,d)\neq(2,3)$. Then for any rank $n$ abelian subgroup $G\subset\mathrm{Aut}(\mathbb{P}^{n+1})$ and any $G$-invariant smooth hypersurface $X\subset\mathbb{P}^{n+1}$ of degree $d$, $X$ is not $G$-birational to an action by a subgroup of the torus on a smooth projective toric variety. In particular, the action is not linearizable.
Proof.
Every rank $n$ abelian group $G$ contains an elementary abelian subgroup $({\mathbb{Z}}/p)^{\oplus n}$, for some prime $p$. If $X$ is not $H$-birational to a toric action for $H\subset G$, it certainly will not be $G$-birational to such an action, so we can and do assume $G\cong({\mathbb{Z}}/p)^{\oplus n}$ from now on.
The group $G\subset\mathrm{PGL}(V)$ has some central extension $\tilde{G}\subset\mathrm{GL}(V)$, where $V\cong k^{n+2}$. We may assume that there is a $G$-fixed point $p$ on $X$, or else $H_{n-1}(\mathcal{D}_{G}(X))=0$, while toric actions satisfy $H_{n-1}\cong{\mathbb{Z}}$. The line $\ell_{p}\subset k^{n+2}$ is a one-dimensional sub-$\tilde{G}$-representation of $V$. After twisting $V$ by this representation (which doesn’t change the image $G$), we can assume that $\tilde{G}$ acts trivially on $\ell_{p}$. It follows that $G$ lifts to $\mathrm{GL}_{n+2}(k)$, and we can diagonalize the action of $G$ on $V$ so that $V\cong\chi_{\mathrm{triv}}\oplus\chi_{1}\oplus\cdots\oplus\chi_{n+1}$. In other words, $\mathbb{P}(V)$ is a compactification of the linear representation $\chi_{1}\oplus\cdots\oplus\chi_{n+1}$ of $G$.
We claim that two of $\chi_{\mathrm{triv}},\chi_{1},\ldots,\chi_{n+1}$ must be the same, as long as $n\geq 3$. (The theorem automatically holds for curves of degree $d\geq 3$ and surfaces of degree $d\geq 4$ because these are all irrational, so we ignore these cases.) Indeed, if the characters are all distinct, the only $G$-fixed points on $\mathbb{P}^{n+1}$ are the coordinate points. We may assume $X$ contains the coordinate point $p$ of $x_{0}$ (which $G$ acts on by the trivial character). Then $T_{p}X\subset T_{p}\mathbb{P}^{n+1}$ is a sub-$G$-representation. Since $T_{p}\mathbb{P}^{n+1}$ is a sum of distinct characters, $T_{p}X$ coincides with the tangent space to a coordinate hyperplane, which we can assume is $\{x_{1}=0\}$ after relabeling. It also follows that $T_{p}X\cong\chi_{2}\oplus\cdots\oplus\chi_{n+1}$ is faithful, so $\chi_{2},\ldots,\chi_{n+1}$ form a basis for $G^{*}\cong\mathbb{F}_{p}^{n}$. By Lemma 2.1, $p$ must be the intersection of $n$ divisors on $X$ which are each fixed by a nontrivial subgroup of $G$. These divisors are precisely $\{x_{2}=0\}\cap X,\ldots,\{x_{n+1}=0\}\cap X$. Indeed, only coordinate strata in $\mathbb{P}^{n+1}$ are fixed by a nontrivial subgroup, and $X$ cannot contain any codimension $2$ coordinate plane of $\mathbb{P}^{n+1}$ when it is smooth of degree $d>1$ and $n\geq 3$ [8, Corollary 6.26].
Since the hyperplanes $\{x_{2}=0\},\ldots,\{x_{n+1}=0\}$ all are fixed by a nontrivial subgroup, the collection $\chi_{1},\ldots,\widehat{\chi_{i}},\ldots,\chi_{n+1}$ does not form a basis of $G^{*}$ for $i=2,\ldots,n+1$. This implies $\chi_{1}$ is trivial, contradicting the assumption that $\chi_{\mathrm{triv}},\chi_{1},\ldots,\chi_{n+1}$ are distinct.
We may therefore suppose that the $G$-action on $V\cong k^{n+2}$ has the following form: each summand of $({\mathbb{Z}}/p)^{\oplus n}$ acts by multiplication by $p$th roots of unity on a distinct coordinate $x_{2},\ldots,x_{n+1}$, and the action on the first two coordinates is trivial. Since $X$ cannot contain a codimension $2$ coordinate plane of $\mathbb{P}^{n+1}$, the only divisors fixed by nontrivial subgroups in $X$ are the intersections $D_{i}\coloneqq X\cap\{x_{i}=0\}$, $i=2,\ldots,{n+1}$, where $D_{i}$ is fixed by the $i$th summand ${\mathbb{Z}}/p$ of $G$. Every intersection of $k<n$ of these divisors has a rank $k$ stabilizer, so it is smooth and irreducible of dimension $n-k\geq 1$. Thus, the $(n-2)$-skeleton of $\mathcal{D}_{G}(X)$ is the same as the $(n-2)$-skeleton of the standard $(n-1)$-simplex. However, $D_{2}\cap\cdots\cap D_{n+1}$ is a union of $d$ distinct points (or else the divisors do not intersect transversely), so $\mathcal{D}_{G}(X)$ has $d$ top-dimensional simplices with identical boundary. Therefore, $H_{n-1}(\mathcal{D}_{G}(X))\cong{\mathbb{Z}}^{d-1}$. By Theorem 2.4 and Example 3.1, $X$ is not $G$-birational to a smooth toric variety with action by a full rank finite subgroup of the torus whenever $d\geq 3$. This implies in particular that the $G$-action on $X$ is not linearizable. ∎
Example 3.3 (Exceptional Cases).
Theorem 3.2 does not hold for the case of quadrics, where $d=2$, and for cubic surfaces, where $(n,d)=(2,3)$. For quadrics, any $G$-action with a $G$-fixed point $p$ is linearizable, becuase the projection from $p$ gives a $G$-equivariant birational map to projective space. Of course, an action by an abelian $G$ without a fixed point is not linearizable, since linearizable actions have fixed points.
There is also a counterexample for smooth cubic surfaces, which are always rational. Let $X=\{x_{0}^{3}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3}=0\}\subset\mathbb{P}^{3}$ be the Fermat cubic surface and $G\subset\mathrm{Aut}(X)$ the normal Klein four-subgroup of the $S_{4}$ permuting the four coordinates. Each involution in $G$ fixes a line and three isolated points in $X$ (these involutions are of type 2B in the classification of Dolgachev-Duncan [7]). No nontrivial element of $G$ fixes a positive genus curve, so $G$ is conjugate to a subgroup of $\mathrm{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$ or $\mathrm{Aut}(\mathbb{P}^{2})$ in the plane Cremona group [3, Theorem 5]. Since $G$ also fixes a point, the subgroup must be linearizable.
Remark 3.4.
There is an alternate approach to the second part of the proof of Theorem 3.2 suggested by Yuri Tschinkel, using the formalism of incompressible symbols in equivariant Burnside groups. We sketch the argument below in the case of $G\coloneqq({\mathbb{Z}}/d)^{\oplus n}$ acting on the first $n$ coordinates of a hypersurface $X$ of dimension $n$ and degree $d$.
[…]
Such a divisor symbol is incompressible roughly if it cannot arise as the symbol for an exceptional divisor of a blowup of a lower dimensional subvariety (for a precise definition, see [12, Definition 3.3]). Assuming by induction the nonlinearizability of the residual action on $k(Y)$, one can prove the incompressibility of the corresponding divisor symbol. It follows that such a symbol cannot appear in the invariant associated to a linearizable action of $G$ on $\mathbb{P}^{n}$.
For every pair $n,d$ in the theorem, there are smooth hypersurfaces to which the result applies. For instance, if
$X\coloneqq\left\{\sum_{i=0}^{n+1}a_{i}x_{i}^{d}=0\right\}\subset\mathbb{P}^{n+1},$
then $X$ is invariant under any subgroup of the action by multiplication by $d$th roots of unity on the coordinates. That group has rank $n+1$.
When the degree $d$ is small compared to the dimension $n$, it is not known whether smooth hypersurfaces $X_{d}\subset\mathbb{P}^{n+1}$ are rational. Even a very general $X$ is only known to be non-rational when the degree exceeds roughly $\log_{2}(n)$ [16]. In contrast, Theorem 3.2 shows that no action by a rank $n$ abelian group on a hypersurface of dimension $n\geq 3$ and degree $d\geq 3$ is linearizable. This suggests the following question:
Question 3.5.
For which finite groups $G\subset\mathrm{Aut}(\mathbb{P}^{n+1})$ and positive integers $n,d$ is it true that the $G$-action on any $G$-invariant smooth hypersurface of degree $d$ in $\mathbb{P}^{n+1}$ is not linearizable?
A positive answer for $G$ trivial and some $n,d$ is equivalent to the statement that no smooth hypersurface of degree $d$ in $\mathbb{P}^{n+1}$ is rational.
Besides the groups in Theorem 3.2, some other interesting cases of 3.5 in low dimensions are known. For instance, if $G\cong{\mathbb{Z}}/3$ is the group acting by third roots of unity on the first coordinate $x_{0}$ in $\mathbb{P}^{3}$, then the induced action on any $G$-invariant smooth cubic surface $X$ is not linearizable, even though $X$ is rational. Indeed, a cyclic group of prime order in the Cremona group $\mathrm{Cr}(2)$ fixing a curve of positive genus is not even stably $G$-birational to a linear action on $\mathbb{P}^{2}$ [4, Corollary 1.2]. However, 3.5 appears to be open for low rank abelian groups acting on hypersurfaces of small degree and large dimension.